OpenStudy (anonymous):

A coffee shop mixes Brazilian coffee worth $5 per kilogram with Turkish coffee worth $8 per kilogram. The mixture is to sell for $7 per kilogram. How much of the Brazilian coffee should be used to make 300 kg of the mixture?

OpenStudy (anonymous):

Well, let's first find out how much of each is needed for one kilogram. What's the ratio needed to make $7 of coffee with the $5 and $8?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Okay, well let's set this up. x=Brazilian coffee So, 300-x=kgs of Turkish coffee x is $5 a kg. So, 5x dollars. 300-x kgs of Turkish coffee is $8 a kg. So it's 8(300-x)dollars Total value of mixture is 5x+8(300-x) dollars We know that it's $7 per kgs Value is 7*300=$2,100 So, our final equation is \[5x+8(300-x)=2100\] Can you figure this out?

OpenStudy (anonymous):

Okay, so we need to find x, right? The whole thing is $2,100. Let's move all of the variables over to one side and the numbers to the other. Can you do this?

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

Go ahead and try :) Or, do you have options to pick from?

OpenStudy (anonymous):

i dont have any options

OpenStudy (anonymous):

i just tried and i still have no idea

OpenStudy (anonymous):

Hmm... I don't know how to explain this to you :-/ Plug in 100 for x and see what you get

OpenStudy (anonymous):

idk i suck at this s***

OpenStudy (anonymous):

Replace every x with 100

OpenStudy (anonymous):

i did but i got a huge a** number wich is the wrong answer